Circular No. 7183

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999br IN NGC 4900 F. Patat, European Southern Observatory (ESO); S. Benetti, Tele-

scopio Nazionale Galileo; and E. Cappellaro, L. Rizzi, and M. Turatto, Padova, write: “We have observed SN 1999br (cf. IAUC 7141) with the ESO 3.6-m telescope (+ EFOSC2) on May 21.02 UT at La Silla. A full reduction of the CCD spectrogram (range 330–1020 nm, resolution 2.0 nm)

shows that this object (http://sc6.sc.eso.org/~fpatat/sn99br.gif) is

remarkably similar to SN 1994N (IAUC 5987) at a comparable phase. The dominant feature is Hα, blended with Ba II 649.7-nm, as shown by the subluminous SN 1997D (Turatto et al. 1998, Ap.J. 498, L129). The other two lines, at 585.4 and 614.2 nm, of the same multiplet 2 of Ba II are visible (the former blended with Na I D). The low expansion velocities allow us to recognize a multitude of other lines, all accompanied by the typical P-Cyg profile. The most prominent features are identified as H, Na I, Ca II, Ba II, Sc II, Fe I, and Fe II. The expansion velocity deduced from the absorption minimum of Hα is 3500 km/s, which is a very low value if compared with ‘normal’ supernovae of type II. In analogy to the more extreme SN 1997D, the small kinetic energy, the low temperature, and the faint magnitude (cf. Filippenko 1999, IAUC 7143) indicate that a small mass of ^{56}Ni has been ejected by SN 1999br. Further observations during the radioactive decay tail are needed to prove this suggestion.”

COMET C/1999 H1 (LEE) E. A. Bergin, Harvard-Smithsonian Center for Astrophysics (CfA);

D. A. Neufeld, Johns Hopkins University; and J. R. Stauffer, S. C. Kleiner, Z. Wang, R. Plume, and G. J. Melnick, CfA, write: “The 1(10)–1(01) transition of water vapor near 557 GHz has been detected toward comet C/1999 H1 by the Submillimeter Wave Astronomy Satellite (SWAS) at a signal-to-noise ratio of 60. Over the period May 19.01–23.69 UT, the average integrated antenna temperature was 1.08 ± 0.02 K km/s within a 3′.3 × 4′.5 (FWHM) elliptical beam. For a Haser model with an assumed water lifetime of 7.3 × 10^{4} s, the inferred production rate for ortho-water is 1.0 × 10^{29} s^{−1}, corresponding to a total water-production rate of 1.3 × 10^{29} s^{−1} if the ortho-to-para ratio is 3. The observed line width of 1.48 ± 0.03 km/s FWHM is broader than the instrumental profile and suggests an intrinsic line width of ∼ 1.0 km/s FWHM. The data, taken during a portion of every 97-min spacecraft orbit over a 4.68-day period, provide no obvious evidence for variability.”

1999 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green